Research shows that 14% of 6th graders in K-6 schools have tried cigarettes and 61% of 7th graders in 7-9 middle school have tried cigarettes.   what is the absolute increase, to the nearest percent?
Answer #1

percentage of students in 6th grade in K-6 schools have tried cigarettes = 14%
percentage of students in 7-9th grade in K-6 schools have tried cigarettes = 61%

Thus, absolute increase = (61-14)= 47%

However,
The percent change (assuming there is 100 students in respective classes) = (61-14)*100/14 = 335.71%

Note: I hope this would help you, but if you still need help in any of the steps then let me know via comments.